# Takin-VC: Zero-shot Voice Conversion via Jointly Hybrid Content and Memory-Augmented Context-Aware Timbre Modeling

### v1

@@ 3.4 Conditional Flow Matching-based Decoder
In Takin-VC, to facilitate more efficient training and faster inference, we leverage a CFM model with optimal-transport (OT-CFM) to approximate the distribution of source Mel-spectrograms and generate outputs conditioned on $x_{s_{ctt}}$ and $x_{t_{cond}}$, all in a simulation-free manner.
Assume that the standard distribution and target distribution are denoted as $p_{0}(x)$ and $p_{1}(x)$, respectively. The OT flow $\phi:[0,1]\times R^{d}\rightarrow R^{d}$ establishes the mapping between two density functions through the use of an ordinary differential equation (ODE):
@@
where $z$ represents the random variable, $\sigma_{min}$ is a hyper-parameter set to 0.0001. As a consequence, the final training objective of the proposed Takin-VC can be formulated as:
@@
where $\theta$ is the weights of the flow matching model, $h$ is the conditional input $x_{t_{cond}}$.
@@ 4.1 Baseline System
We conduct a comparative experiment of the performance in zero-shot voice conversion between our proposed Takin-VC approach and baseline systems, encompassing the following system:
  • DiffVC [39]: A zero-shot VC system based on diffusion probabilistic modeling, which employs an averaged mel spectrogram aligned with phoneme to disentangle linguistic content and timbre information.
  • NS2VC [^5] https://github.com/adelacvg/NS2VC: A modified voice conversion edition of NaturalSpeech2 [64], which employ both diffusion and codec model to achieve zero-shot VC.
  • VALLE-VC [65]: We replace the original phoneme input with the semantic token extracted from the supervised model to make VALLE convert the timbre of source speech to the target speaker.
  • SEFVC [8]: A speaker embedding free voice conversion model, which is designed to learn and incorporate speaker timbre from reference speech via a powerful position-agnostic CA mechanism and then reconstruct waveform from HuBERT semantic tokens in a non-autoregressive manner.
@@ 4.2 Evaluation Metrics
To evaluate the performance of our proposed Takin-VC and baseline systems, both subjective and objective metrics are introduced. For subjective metrics, we employ naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples and similarity mean opinion scores (SMOS) to evaluate the speaker similarity. We invite 20 professional participants to listen to the generated samples and provide their subjective perception scores on a 5-point scale: ’5’ for excellent, ’4’ for good, ’3’ for fair, ’2’ for poor, and ’1’ for bad. For objective metrics, we employ word error rate (WER), UTMOS, and speaker embedding cosine similarity (SECS) to evaluate the intelligibility, quality, and speaker similarity. Specifically: 1) We use a pre-trained CTC-based ASR model [^6] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe the generated speech and compare with ground-truth transcription; 2) We use a MOS prediction system that ranked first in the VoiceMOS Challenge 2022777https://github.com/tarepan/SpeechMOS to estimate the speech quality of the generated samples; 3) We use the WavLM-TDCNN speaker verification model [^8] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification to measure speaker similarity between generated speech and target speech.
@@ 4.4 Model Configuration
For the content encoder part, in the first stage, we used the 12-layer HybridFormer-base model trained on a large dataset of 500K hours. For the wavlm part, we used the output features of the 6th layer. In the VQ part, we adopted a single-layer 8200 codebook with a hidden dimension of 1024, trained for 1 million steps on 100K hours of data. The fusion layer, as described in Sec. 3.2, is a simple process of conv1d, activation layer, and weighted summation. The Decoder adopts the same structure and configuration as Hificodec [44].
In the part of timbre modeling and flow matching restoration, both the context-aware timbre module and the memory module use a transformer layer with 8 heads, 6 layers, and a hidden size of 1024, with only the form of attention being different. The main structure of flow matching uses a design of 10-layer Unet plus 3 layers of resblock, with a hidden size of 1280. A Memory Fusion Block is inserted into the 10-layer Unet to enhance the timbral similarity of the generated audio.
For the small-data experiments, we used four A800 GPUs, whereas the large-data experiments were conducted on eight A800 servers. The batch size on each card was set to 16, and the AdamW learning rate was set to 1e-4. In the inference section, experiments typically took 15 to 50 steps, with the final table uniformly adopting the results of 50 steps. The Classifier-Free Guidance (CFG) coefficient ranged from 0.1 to 1.0, with 0.7 used in the table. The specific experimental results will be detailed later.
@@ 5.1 Experiments on small dataset
GT 4.17±$\pm 0.04$ 2.04 4.21 -
DiffVC 3.75±$\pm 0.05$ 3.66±$\pm 0.07$ 3.08 3.68 0.61
NS2VC 3.65±$\pm 0.07$ 3.51±$\pm 0.06$ 2.94 3.64 0.53
VALLE-VC 3.80±$\pm 0.06$ 3.79±$\pm 0.04$ 2.77 3.72 0.65
SEFVC 3.68±$\pm 0.05$ 3.76±$\pm 0.06$ 3.75 3.51 0.63
Takin-VC 3.98±$\pm 0.04$ 4.11±$\pm 0.05$ 2.35 4.08 0.71
Furthermore, we evaluate the performance using objective metrics. The WER of our proposed system is 2.35, only slightly higher than the ground truth samples, indicating that the samples generated by Takin-VC exhibit better intelligibility. Moreover, Takin-VC achieves a UTMOS of 4.08 and an SECS of 0.71, demonstrating superior quality and similarity performance. Overall, the objective results of our proposed Takin-VC outperform all baseline systems and further corroborate the subjective findings.
@@ 5.3 Ablation Study
We conduct ablation experiments to evaluate the effectiveness of each component in our proposed system for generating natural-sounding samples and accurate timbre modeling.
Table 3: Experimental results on ablation studies. “w/o vp” represents removing voice print in the attention module. “w/o hybrid” represents replacing the proposed hybrid content encoder with the conventional used ASR encoder, and “w/o memory” means removing the timbre memory module.
@@
   w/o hybrid 3.67±$\pm 0.04$ 4.01±$\pm 0.04$ 2.79 3.89 0.66
@@
   w/o memory 3.92±$\pm 0.04$ 3.75±$\pm 0.05$ 2.44 4.01 0.52
As shown in Table 3, NMOS and WER results degrade when we replace the proposed hybrid content encoder with a conventional ASR encoder. This suggests that the conventional ASR encoder is less capable of disentangling linguistic content from the necessary paralinguistic information, underscoring the importance and effectiveness of our hybrid encoder in extracting linguistic content. Additionally, we observe a notable decline in speaker similarity when the voice print is removed from the attention module. We believe the voice print introduces a stronger timbre bias, which helps the attention module focus on capturing timbre information. Furthermore, when we remove the memory module, SMOS and SECS scores show significant degradation compared to the original Takin-VC, demonstrating the critical role of the memory module in improving timbre modeling. These ablation results demonstrate the effectiveness of each component proposed in our Takin-VC.
6 Discussion and Limitations
Takin is an effective and data-efficient zero-shot VC system that achieves comparable naturalness and speaker adaptation performance to its large-scale, autoregressive counterparts. The core of this approach lies in the neural codec training based hybrid linguistic content encoder, which captures high-quality speaker-agnostic content representations, and the introduction of both context-aware timbre modeling and memory-augmented modules to enhance speaker similarity performance. In many ways, our work provides a strong foundation for future studies, as we demonstrate that state-of-the-art performance in this task can be achieved without relying on complex training setups, representation quantization steps, or costly autoregressive models.
This work primarily focuses on zero-shot capabilities for speech generation, while zero-shot capabilities for speech editing remain limited and are a subject for future exploration. Additionally, while high-quality zero-shot VC has great potential, it can also lead to negative social impacts, such as voice impersonation of public figures and non-consenting individuals. We highlight this as a potential misuse of the technology to raise awareness of its ethical implications.
7 CONCLUSIONS
In this study, we propose a novel framework called Takin-VC, designed to achieve high quality and similarity in zero-shot VC. We introduce an effective neural codec training guided hybrid content encoder that leverages quantized features from both pre-trained HybridFormer and WavLM to extract the linguistic content of the source speech. This hybrid content encoder improves the naturalness and intelligibility of the converted speech. Additionally, we present an advanced cross-attention-based, context-aware timbre modeling approach that captures fine-grained, semantically associated target timbre features. Furthermore, we employ a conditional flow-matching model to efficiently reconstruct the Mel-spectrogram of the source speech and propose an efficient memory-augmented module for the flow-matching process, enhancing the overall performance of the generated samples. Experimental results demonstrate that Takin-VC outperforms all baseline systems in naturalness and speaker similarity on benchmark datasets. Ablation studies also confirm the effectiveness of each component in our approach.

### v2

@@ 3.4 Conditional Flow Matching-based Decoder
In Takin-VC, to facilitate more efficient training and faster inference, we use a CFM model with optimal-transport (OT-CFM) to approximate the distribution of source Mel-spectrograms and generate predicted outputs conditioned on $X_{s_{ctt}}$ and $X_{t_{tcond}}$, all in a simulation-free manner.
Assume that the standard distribution and the target distribution are denoted as $p_{0}(x)$ and $p_{1}(x)$, respectively. The OT flow $\phi:[0,1]\times R^{d}\rightarrow R^{d}$ establishes the mapping between two density functions through the use of an ordinary differential equation (ODE):
@@
where $z$ represents the random variable, $\sigma_{min}$ is a hyper-parameter set to 0.0001. Therefore, the training objective of the proposed CFM model can be formulated as:
@@
where $\theta$ represents the parameters of the flow matching model, and $h$ denotes the conditional set comprising $X_{t_{tcond}}$ and $X_{s_{c}t_{t}}$.
3.5 Training Objective
The training objective of the proposed Takin-VC is composed of two components, i.e., the RVQ commitment loss $\mathcal{L}_{vq}$ of the VQ module and $\mathcal{L}_{cfm}$.
$\begin{gathered}\mathcal{L}_{total}=\mathcal{L}_{cfm}+\lambda\mathcal{L}_{vq}\\ \mathcal{L}_{vq}(X_{s},\tilde{X_{s}})=\sum_{i=1}^{N}\left\|X_{s_{i}}-\hat{X}_{s_{i}}\right\|_{2}^{2}\end{gathered}$ (5)
Here, $\lambda$ is a hyper-parameter that controls the weight of $L_{vq}$, and $N$ represents the number of RVQ-based quantizers. In our implementation, $\lambda$ is empirically set to 0.01, $N$ is set to 1, and the codebook size of the RVQ-based quantizer is empirically determined to be 8200.
@@ 4.1 Baseline System
We conduct a comparative experiment of the performance in zero-shot voice conversion between our proposed Takin-VC approach and baseline systems, encompassing the following system: 1) DiffVC [36]: A zero-shot VC system based on diffusion probabilistic modeling, which employs an averaged mel spectrogram aligned with phoneme to disentangle linguistic content and timbre information; 2) NS2VC [^5] https://github.com/adelacvg/NS2VC: A modified voice conversion edition of NaturalSpeech2 [57], which employ both diffusion and codec model to achieve zero-shot VC; 3) VALLE-VC [58]: We replace the original phoneme input with the semantic token extracted from the supervised model to make VALLE convert the timbre of source speech to the target speaker; 4) SEFVC [34]: A speaker embedding free voice conversion model, which is designed to learn and incorporate speaker timbre from reference speech. 5) StableVC [59]: A style controllable zero-shot voice conversion system, which employs dual adaptive gate attention to capture timbre and style information. 6) SeedVC [60]: A zero-shot voice conversion system with an external timbre shifter and diffusion transformer.
@@ 4.2 Evaluation Metrics
Both subjective and objective metrics are employed to evaluate the performance of our Takin-VC and baseline systems. For subjective metrics, we employ naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples and similarity mean opinion scores (SMOS) to evaluate the speaker similarity. We invite 20 professional participants to listen to the generated samples and provide their subjective perception scores on a 5-point scale: ’5’ for excellent, ’4’ for good, ’3’ for fair, ’2’ for poor, and ’1’ for bad. For objective metrics, we employ word error rate (WER), UTMOS, and speaker embedding cosine similarity (SECS) to evaluate the intelligibility, quality, and speaker similarity. Specifically: 1) We use a pre-trained CTC-based ASR model [^6] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe the generated speech and compare with ground-truth transcription; 2) We use a MOS prediction system that ranked first in the VoiceMOS Challenge 2022777https://github.com/tarepan/SpeechMOS to estimate the speech quality of the generated samples; 3) We use the WavLM-TDCNN SV model [^8] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification to measure speaker similarity between generated speech and target speech. Furthermore, we introduce real-time factor (RTF) to evaluate the efficiency of Takin-VC.
@@ 4.4 Model Configuration
For the content encoder part, in the first stage, we used the 12-layer HybridFormer-base model trained on a large dataset of 500K hours. For the WavLM part, we used the output features of the 6th layer. In the VQ part, we adopted a single-layer 8200 codebook with a hidden dimension of 1024, trained for 1 million steps on 100K hours of data. The fusion layer, as described in Sec. 3.2, is a simple module with several convolutional layers, an activation layer, and weighted summation. The Decoder adopts the same structure and configuration as HiFi-codec [41].
In the part of timbre modeling and flow matching model, both the context-aware and memory-augmented modules use a transformer block with 8 heads, 6 layers, and a hidden size of 1024, with only the form of attention being different. The main structure of CFM uses a design of 10-layer U-net plus 3 layers of ResNet block [61], with a hidden size of 1280. A Memory Fusion Block is inserted into the 10-layer U-net to enhance the speaker similarity of the generated audio.
For the small-data experiments, we use four A800 GPUs, whereas the large-data experiments are conducted on eight A800 servers. The batch size on each GPU is set to 16 with the AdamW optimizer using 1e-4 as the learning rate. In the inference section, experiments typically took 5 to 20 steps, with the final table uniformly adopting the results of 10 steps. The Classifier-Free Guidance (CFG) coefficient ranged from 0.1 to 1.0, with 0.7 used in the table. The specific experimental results will be detailed later.
@@ 5.1 Experiments on small dataset
NMOS ($\uparrow$) SMOS ($\uparrow$) WER ($\downarrow$) UTMOS ($\uparrow$) SECS ($\uparrow$) RTF ($\downarrow$)
GT 4.17±$\pm 0.04$ 2.04 4.21 -
DiffVC 3.75±$\pm 0.05$ 3.66±$\pm 0.07$ 3.08 3.68 0.61 0.294
NS2VC 3.65±$\pm 0.07$ 3.51±$\pm 0.06$ 2.94 3.64 0.53 0.347
VALLE-VC 3.80±$\pm 0.06$ 3.79±$\pm 0.04$ 2.77 3.72 0.65 3.678
SEFVC 3.68±$\pm 0.05$ 3.76±$\pm 0.06$ 3.75 3.51 0.63 0.187
StableVC 3.83±$\pm 0.04$ 3.88±$\pm 0.06$ 2.77 3.92 0.66 0.267
SeedVC 3.87±$\pm 0.05$ 3.74±$\pm 0.06$ 2.51 3.81 0.68 0.341
Takin-VC 3.98±$\pm 0.04$ 4.11±$\pm 0.05$ 2.35 4.08 0.71 0.154
@@
Furthermore, we evaluate the performance using objective metrics. The WER of our proposed system is 2.35, only slightly higher than the ground truth samples, indicating that the samples generated by Takin-VC exhibit better intelligibility. Moreover, Takin-VC achieves a UTMOS of 4.08 and an SECS of 0.71, demonstrating superior quality and similarity performance. Overall, the objective results of our proposed Takin-VC outperform all baseline systems and further corroborate the subjective findings. For inference efficiency, Takin-VC achieves the lowest RTF over all baseline systems, demonstrates superior real-time performance.
@@ 5.3 Ablation Study
Table 3: The ablation results for linguistic content extraction modules. “w/o ppg” and “w/o SSL” represent removing the HybridFormer or WavLM branch in our proposed hybrid content encoder, respectively.
NMOS SMOS WER UTMOS SECS
@@
   w/o ppg 3.74±$\pm 0.04$ 3.07±$\pm 0.04$ 2.79 3.91 0.45
   w/o SSL 3.63±$\pm 0.04$ 3.81±$\pm 0.04$ 2.64 3.84 0.67
We conduct two ablation experiments to evaluate the effectiveness of each proposed component in linguistic content extraction and timbre modeling. As shown in Table 3, SMOS results are significantly degraded, suggesting that only using the SSL model to extract linguistic content will result in timbre leakage. When we remove the SSL model in the hybrid content encoder and only use HybridFormer to extract linguistic content, we can find that NMOS and WER results degrade. This suggests that the conventional ASR encoder is less capable of disentangling linguistic content from the necessary paralinguistic information, underscoring the importance and effectiveness of our hybrid encoder in extracting linguistic content.
Table 4: The ablation results for timbre-related modules. “w/o con” represents removing content-aware timbre modeling and only employing voice print to extract timbre information. “w/o vp” represents removing the voice print, and “w/o mem” means removing the memory-augmented timbre modeling module.
NMOS SMOS WER UTMOS SECS
Takin-VC 3.98±$\pm 0.04$ 4.11±$\pm 0.05$ 2.35 4.08 0.71
   w/o con 3.77±$\pm 0.04$ 3.61±$\pm 0.04$ 3.01 3.85 0.58
@@
   w/o mem 3.92±$\pm 0.04$ 3.75±$\pm 0.05$ 2.44 4.01 0.52
Additionally, we conduct an ablation study for timbre-related modules, results are shown in Table 4. We find significant degradation across all metrics when removing context-aware timbre modeling. It suggests that the system can not capture timbre information as well without the module, resulting in poor generation results. We observe a notable decline in speaker similarity when the voice print is removed from the attention module. We believe the voice print introduces a stronger timbre bias, which helps the attention module focus on capturing timbre information. Furthermore, when we remove the memory-augmented timbre modeling module, SMOS and SECS scores show significant degradation compared to the original Takin-VC, demonstrating the critical role of the memory module in improving timbre modeling. These ablation results demonstrate the effectiveness of each component proposed in our Takin-VC.
6 Conclusion
In this study, we introduce Takin-VC, an effective framework for expressive zero-shot VC. Leveraging an adaptive fusion-based hybrid content encoder, Takin-VC integrates the complementary strengths of PPGs and quantized WavLM features in a learnable manner, thereby enhancing the naturalness and expressiveness of the converted speech. To improve speaker similarity, we propose an advanced memory-augmented module capable of extracting fine-grained conditional target timbre features. Additionally, we design a context-aware timbre modeling module to capture fused representations that effectively align and exploit the source content with target timbre elements. To enable stable training and fast inference, a conditional flow-matching model is presented reconstruct the Mel-spectrogram of the source speech. Experimental results demonstrate that Takin-VC outperforms all baseline systems regarding naturalness, expressiveness, speaker similarity, and real-time performance. Ablation studies further validate the effectiveness of each proposed component in our framework.